OpenStudy (anonymous):

5. -7x^2+6x+3=0 and -8x^2-8x-2=0 for each one how many real number soloutions does the equation have?

OpenStudy (anonymous):

use discriminant: b^2-4ac in quadratic equation ax^2+bx+c=0 if discriminant > 0, 2 solutions if =, 1 solution if < 0, no solutions

OpenStudy (anonymous):

I dont understand

OpenStudy (anonymous):

so for the first one the discriminant is (6)^2-4(-7)(3) = 36+84 = 120 Therefore, two solutions

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

\[\begin{array}{l}{x_{1/2}} = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\\\\\Delta = {b^2} - 4ac\\\Delta > 0\\\Delta = 0\\\Delta < 0\end{array}\]

OpenStudy (anonymous):

I can explain further if necessary

OpenStudy (anonymous):

please do

OpenStudy (anonymous):

OK then...

OpenStudy (anonymous):

Here's the graph of y = -7x^2+6x+3 From the graph we can see that -7x^2+6x+3 = 0 is true at 2 points Therefore it has 2 solutions

OpenStudy (anonymous):

what about the second one?

OpenStudy (anonymous):

-8x^2-8x-2=0 Let's graph it

OpenStudy (anonymous):

\[\LARGE -7x^2+6x+3=0\] now use the quadratic formula: \[\LARGE {x_{1/2}} = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\] \[\LARGE \begin{array}{l}{x_{1/2}} = \frac{{ - 6 \pm \sqrt {36 - 4( - 7)(3)} }}{{2( - 7)}} \ \ \\ \\ \\{x_{1/2}} = \frac{{ - 6 \pm \sqrt {36 + 84} }}{{ - 14}}\\{x_{1/2}} = \frac{{ - 6 \pm \sqrt {120} }}{{ - 14}}\end{array}\] so...? (the first one)

OpenStudy (anonymous):

We can see it only touches the x axis (where y = 0) once, so it has one solution. Using the discriminant: (-8)^2 - 4(-8)(-2) = 0. Therefore only one solution

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

No probs